### v2

# Meta-GGAs vs. Hybrid Functionals for Point Defects: The Best of Both Worlds Applied to Layered MnO2, NiO2 and KCoO2

Abstract
Defects in a material can significantly tune properties and enhance utility. Hybrid functionals like HSE06 are often used to describe solids with such defects. However, geometry optimization (including accounting for effects such as Jahn-Teller distortion) using hybrid functionals is challenging for the large supercells needed for defect study. The proposed r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method, which is computationally much cheaper and faster than hybrid functionals, can successfully describe defects in materials with the proper choice of U (for the d orbitals of the host atom) and $\mathrm{U_{d}}$ (for those of the defect atom), as shown here for small polaron defects in layered transition-metal oxides. We use a literature value of U or $\mathrm{U_{d}}$ appropriate to a given transition-metal ion and its oxidation state. The layered materials birnessite ($\mathrm{K_{n}MnO_{2}},n=0.03$) and $\mathrm{K_{n}NiO_{2}},n=0.03$, with one K atom intercalated between layers in a supercell, are found to have one localized occupied $\mathrm{e_{g}}$ polaronic state on the transition metal ion reduced by the insertion of the K atom, when the geometry is calculated as above using published U values. The expected Jahn-Teller distortion is not observed when U=$\mathrm{U_{d}}$=0. Layered cobalt oxide with additional potassium ions intercalated ($\mathrm{K_{n}CoO_{2}},n=1.03$) is different, due to a dramatic difference in electronic configuration of the defected Co(II) ion: A single extra K atom in the supercell leads to four localized electrons in the band gap, using standard U values, and even for U=$\mathrm{U_{d}}$=0.
I Introduction
This work is part of a larger project on the computational identification of alkali-atom-intercalated layered materials as promising catalysts for the oxygen evolution reaction (OER) of water splitting for clean hydrogen production. Earlier work Peng et al. (2017); Ding et al. (2021) showed the importance of small-polaron defects formed by transfer of an electron from an intercalated alkali atom to a neighboring manganese ion. Calculation of the polaron often requires a nonlocal density functional for the exchange-correlation energy, and geometry optimization in a large supercell. We found that hybrid functionals as used in Ref. Peng et al. (2017) are too expensive for a broad materials search. In this article, we show that an alternative r2SCAN+rVV10+U+$\mathrm{U_{d}}$ approach can achieve comparable accuracy at much lower cost, and may be useful for other point defects in a range of materials. We also found that, while one intercalated K atom creates one localized electron in the energy gap of layered MnO2 and NiO2, it can create four in layered KCoO2. We further found that, while r2SCAN needs a +U correction to create a polaron in MnO2 and NiO2, it does not need one in KCoO2. Our broader computational and experimental materials search is now underway.
Various types of defects exist in solids, and defects in solids can influence numerous important properties like electrical conductivity, reactivity, and magnetic or optical properties. For example, the Oxygen Evolution Reaction (OER) is frequently favored by defects like polarons in transition metal oxide (TMO)-based catalysts Peng et al. (2017); Ding et al. (2021). Leveraging defects as a tool allows fine tuning of the electronic properties of materials, making the understanding of defects in materials a pivotal area of research.
The computational design and study of such materials using first-principles density functional theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965) offer valuable early insights. However, the approach presents challenges, as density functional approximations (DFAs) introduce self-interaction error (SIE). Popular DFAs such as LDA/GGA/meta-GGA tend to underestimate the Perdew-Parr-Levy-Balduz (PPLB) Perdew et al. (1982) straight line condition, leading to inaccuracies in describing charge transfer that are reduced but not eliminated in that sequence of functionals. Given that defect studies require proper charge transfer, DFAs often fall short in accurately portraying defects in a system.
Hybrid functionals, which combine a fraction of exact exchange, such as HSE06 Krukau et al. (2006a) which utilizes 25% exact exchange and 75% PBE exchange in the short range along with full PBE exchange in the long range, experience less SIE and often provide a more accurate description of the electronic structure of materials. Hybrid functionals have been successfully employed in studying defects in solids. For instance, Peng et al. Peng et al. (2017) effectively investigated polaron-like defects in birnessite (Fig. 1, $\mathrm{K_{n}MnO_{2}},n<1$). Nevertheless, the inclusion of exact exchange in hybrid functionals renders them computationally expensive. The structural relaxation using hybrid functionals becomes particularly costly for a reasonably sized supercell with localized defects, with computational expenses rapidly escalating as the supercell size increases toward better simulation of defects. Consequently, the structural optimization of large supercells becomes impractical. Another challenge associated with hybrid functionals is that a material-independent exact exchange mixing parameter is not determined through any exact condition, nor is the range-separation parameter in range-separated hybrids Sun et al. (2016). Peng et al. determined a mixing parameter of 0.22 to study defects in birnessite using HSE06 Peng et al. (2017), deviating from the original value of 0.25. Additionally, determining a mixing parameter for semiconductors that may not be suitable to metals or insulators Henderson et al. (2011) adds to the challenge of finding an appropriate parameter for the system under study when employing hybrid functionals.
$\mathrm{r^{2}SCAN}$ Furness et al. (2020a) is a recently developed meta-generalized gradient approximation that reinstates exact constraint adherence to rSCAN Bartók and Yates (2019), preserving the numerical efficiency of rSCAN while simultaneously restoring the transferable accuracy of SCAN Furness et al. (2020b). Several studies have demonstrated that SCAN predicts geometries and other properties as well as or even better than hybrid functionals. Sun et al. Sun et al. (2016) showed that SCAN accurately predicts geometries and energies of diversely bonded materials and molecules, matching or surpassing the accuracy of computationally expensive hybrid functionals. Another study by Saÿnick and Cocchi Saÿnick and Cocchi (2021) on cesium-based photocathode materials $\mathrm{Cs_{3}Sb}$ and $\mathrm{Cs_{2}Te}$ reported excellent performance of SCAN and HSE06 for both structural and electronic properties. A recent paper on the arXiv Wickramaratne and Lyons reports that, while SCAN may not reliably describe the properties of deep defects and small polarons in several semiconductors and insulators, it yields remarkably good agreement with experimental structural parameters for materials like ZnO, GaN, $\mathrm{Ga_{2}O_{3}}$ and NaF. Additionally, a study by Varadwaj and Miyake Varadwaj and Miyake (2022) on the geometrical, electronic, and optical properties of vanadium dioxide found that SCAN and SCAN-rVV10 can adequately predict the most important geometrical and optoelectronic properties of $\mathrm{VO_{2}}$. Numerous related studies further support the idea that SCAN successfully describes the structure and other properties of materials. Given that $\mathrm{r^{2}SCAN}$ closely agrees with SCAN in accuracy, we expect that $\mathrm{r^{2}SCAN}$ would exhibit similar accuracy in the aforementioned studies.
By construction, $\mathrm{r^{2}SCAN}$ can exhibit very small or negligible SIE (as reflected by its smaller Hubbard U correction) and performs at the level of hybrid functionals but demands less computational resource and time. In this study, we demonstrate that the $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$ functional can effectively describe the defects in materials at the level of the hybrid functional HSE06, or possibly even better. Here, U represents the Hubbard U correction of Anisimov and collaborators Anisimov et al. (1991, 1993); Solovyev et al. (1994), applied to transition metal sites other than the defect site, while $\mathrm{U_{d}}$ is the correction applied to the defect site. Cococcioni and de Gironcoli Cococcioni and de Gironcoli (2005) showed that the +U correction can be regarded as a many-electron self-interaction correction that, like the PPLB condition Perdew et al. (1982), penalizes non-integer electron number on a localized orbital to which it is applied.
Additionally, vdW denotes the long-range van der Waal’s correction. In our approach, we utilize rVV10 Vydrov and Voorhis (2010); Sabatini et al. (2013) to account for this interaction, which importantly reduces inter-layer spacing. In many of our calculations, including HSE06 without vdW, we use the supercell volume and shape from $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$, although we can as a check relax the internal coordinates with HSE06.
[…]
However, this situation is not universal, as we can see that U values for Mn for $\mathrm{r^{2}SCAN}$ are 1.5 eV, 2.1 eV and 1.8 eV for Mn ions with average oxidation states +3.5, +2.5 and +2.33 respectively Swathilakshmi et al. (2023). Here we see that Mn in the +2.5 average OS needs more U correction than Mn in the +2.33 average OS, which does not follow the trend observed for vanadium, although U for Mn in the +3.5 average OS supports the trend within the same elemental series. However it is obvious for the Mn oxide system that Mn ions in different oxidation states need different U corrections. Another example for Co is found in Ref. Long et al. (2020), which found U=3.0 eV as a correction to SCAN for the oxidation reaction 6CoO+$\mathrm{O_{2}\rightarrow 2Co_{3}O_{4}}$ (in which the Co ions are in the +2 and +2.67 oxidation states), but also found that SCAN with this U wrongly predicted the meta-stable crystal O1-CoO2 (in which the Co oxidation state is +4) to be non-metallic, while SCAN with U=0 correctly predicted it to be a metal. Thus the U value can vary with the oxidation state of ions in the system. This indicates that if a system has transition metal ions in different OS, the proper way to describe such a system would be to apply different U corrections to the transition metal ions based on their OS. This could be because DFAs make different SIE for ions in different OS.
Here we present a study of defects in solids by the double U correction $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$ method, where $\mathrm{U_{d}}$ is the U correction applied to the defect site and U is the U correction applied to remaining sites as required for the system. Since U typically increases with decreasing OS, and defect sites are typically in a lower oxidation state, we expect $\mathrm{U_{d}\geq U}$, although exceptions are possible.
One can use the $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$ geometry and the HSE06 hybrid functional orbital energies to leverage the best features of the hybrid functional (e.g., band gap). It appears that employing the full $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$ and/or HSE06 using $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$ geometry is a state-of-the-art method for efficient study of defects in solids.
II Results and Discussion
We choose three layered pristine TMOs, $\mathrm{MnO_{2}}$, $\mathrm{NiO_{2}}$, and $\mathrm{KCoO_{2}}$ as the starting point for our study. Transition metal and oxygen ions in these TMOs are arranged in layers of MO6 (where M=Ni, Mn, Co) octahedra, with 32 transition metal ions per supercell. Inserting an additional potassium ion into the supercell between the layers creates a polaronic defect, specifically a Jahn-Teller electron small polaron Peng et al. (2017). This defect has been studied in birnessite ($\mathrm{MnO_{2}}$) using the HSE06 functional Peng et al. (2017).
Figure 1: General structure for layered potassium metal oxides $\mathrm{K_{n}MnO_{2}}$.
II.1 Layered MnO2
[…]
Figure 2: HSE06 ($\alpha$ = 0.22) spin-resolved density of states per atom in (A) pristine MnO2 with Mn(IV) ions and (B) MnO2 with a single intercalated K+ ion and a defect Mn(III) cation in a supercell with 97 atoms. For comparison with Ref. Peng et al. (2017), the interlayer spacing is set to 7.12 Å, and internal coordinates are relaxed in HSE06. In Figs. 1-9 (except in Figs. 5, and 8), the shaded area shows the total density of one-electron states, and the blue curve shows the transition-metal-d states projected onto the site where the small polaron forms or is expected to form. Also in Figs. 1-9 (except in Figs. 5, and 8), the peaks (if any) in the band gap below the conduction band minimum (CBM) are occupied, localized states of the defect ion. The energy zero is set to the energy of the lowest-energy unoccupied orbital. The total density has been scaled down to make it comparable in size to the projected Mn-d density of states.
Figure 3: r2SCAN+rVV10+U+$\mathrm{U_{d}}$ spin-resolved density of states per atom in (A) pristine MnO2 with U=$\mathrm{U_{d}}$=1.8 eV and in a single K-intercalated MnO2 with (B) U = $\mathrm{U_{d}}$ = 0.0 eV, (C) U= $\mathrm{U_{d}}$ =1.8 eV, (D) U=1.71 eV and $\mathrm{U_{d}}$ =1.58 eV. The geometry has been optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing  5.26 Å. The insets in Figs. A and B show the DOS near the 0 eV region. The inset of Fig. B shows that for U=$\mathrm{U_{d}}$=0 the extra electron from the intercalated K goes to the bottom of the conduction band, making a semi-metal.
We will see similar effects upon a K-intercalation in $\mathrm{NiO_{2}}$ and $\mathrm{KCoO_{2}}$ systems later, and these effects can be understood through similar reasoning. The d-state electronic configurations of Mn(IV) in pristine $\mathrm{MnO_{2}}$ and Mn(III) in a single-K-intercalated $\mathrm{MnO_{2}}$ are as shown in Fig. 5 A and B respectively.
Figure 4: HSE06($\alpha$=0.25)+D3 spin-resolved density of states per atom using r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry of Fig. 3 in (A) pristine MnO2 with U=$\mathrm{U_{d}}$=1.8 eV and in a single K-intercalated MnO2 with (B) U=$\mathrm{U_{d}}$=0.0 eV, (C) U=$\mathrm{U_{d}}$=1.8 eV, (D) U=1.71 eV and $\mathrm{U_{d}}$=1.58 eV. The insets in Figs. A and B show DOS near the 0 eV region. The inset in Fig. B shows that for U=$\mathrm{U_{d}}$=0 the extra electron from the intercalated K goes to the bottom of the conduction band, making a semi-metal.
Figure 5: d-orbital splitting (A) in Mn(IV) in pristine MnO2 demonstrating filled t2g states and empty eg states and (B) in Mn(III) in a single K-intercalated MnO2.
[…]
First, we performed $\mathrm{r^{2}SCAN+rVV10+U}$ calculations for pristine birnessite with the optimal U of 1.8 eV and obtained the DOS as shown in Fig. 3A. $\mathrm{r^{2}SCAN+rVV10+U}$ predicts a band gap of 2.30 eV for the pristine structure. The underestimation of the band gap compared to HSE06 by meta-GGA/GGA is a well-known general trend Brothers et al. (2008). Next, we performed $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$ calculations for a K-intercalated MnO2. As $\mathrm{r^{2}SCAN}$ exhibits a small SIE, we initially set U=$\mathrm{U_{d}}$=0.0 eV to see the performance of r2SCAN+rVV10 without U correction. However, this method failed to resolve the Jahn-Teller defect, as depicted in Fig. 3B. The corresponding electron from the K atom is delocalized, which can be seen by the extra electrons between the CB minimum and the chemical potential or Fermi level at 0 eV. This can be clearly seen in the inset of Fig. 3B. For the description of a polaron in a single K-intercalated MnO2, DFA has to transfer charge from the inserted K atom to a Mn atom. The charge delocalization error of DFAs prohibits transferring a complete electron from the K atom to the Mn site. This indicates the need for +U correction, which can remove the partial occupancy and localize the electron on the defect site. Subsequently, we tested the optimal U by setting U=$\mathrm{U_{d}}$=1.8 eV and the obtained DOS is shown in Fig. 3C, revealing the appearance of a polaronic peak, and the appropriate distortion of Mn-O bonds at the Jahn-Teller distorted manganese atom (see Fig. S2 B). The +U correction applies an energy penalty to the partially occupied orbital. As a result the polaron is localized and appears at the CB edge. We see shifting of Mn(III) d-states to higher energy, and spin symmetry breaking in the total DOS in the CB, consistent with the HSE06 results. As discussed in the work by Peng et al. Peng et al. (2017) and Ding et al. Ding et al. (2021), polaron formation can with the right distribution of intercalated potassium atoms create a potential step between layers that facilitates electron transfer between layers and enhances catalytic activity.
Motivated by studies indicating that U depends on OS, and that generally U decreases with an increase in OS, we determined U for the r2SCAN+rVV10+U functional for Mn(III) and Mn(IV) by comparing r2SCAN+rVV10+U’s magnetic moment with the HSE06+D3 magnetic moment. For this we used pristine KMnO2 and MnO2 systems. First, we relaxed the structure using HSE06+D3 and used the structure to get the magnetic moment due to spin charge density within the Wigner-Seitz sphere around an Mn ion. We used the HSE06+D3 geometry for r2SCAN+rVV10+U and slowly increased U in small steps until we obtained the same magnetic moment as HSE06+D3. This method gave U values of 1.71 eV and 1.58 eV for Mn(III) and Mn(IV) ions, respectively. These values are close to the values in Ref. Swathilakshmi et al. (2023).
We then obtained the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ DOS of a K intercalated MnO2 with U=1.71 eV and $\mathrm{U_{d}}$=1.58 eV which is as shown in Fig. 3C. U=1.71 eV and $\mathrm{U_{d}}$=1.58 eV gave similar results to U=$\mathrm{U_{d}}$=1.8 eV.
We also performed HSE06+D3 calculations on the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometries, and the resulting DOS are plotted in Fig. 4. The D3 dispersion correction was used for the HSE06 functional with new D3 damping parameters a1 = 0.383, a2 = 5.685, and s8 = 2.310 generated for the HSE06 functional Jonas and Grimme (2014). For HSE06+D3 calculations, we used an exact exchange mixing parameter ($\alpha$) of 0.25. This is because we found that the r2SCAN+rVV10+U geometry in much better agreement with the HSE06+D3 geometry for $\alpha$=0.25 than for 0.22. More details are in the Supplementary Material.
Fig. 4A shows the HSE06+D3 DOS of pristine MnO2 obtained using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (U=$\mathrm{U_{d}}$=1.8 eV) geometry. HSE06+D3 increases the band gap of pristine MnO2, close to the HSE06 result (Fig. 2A). Fig. 4B shows the HSE06+D3 DOS of a K-intercalated MnO2 obtained using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (U=$\mathrm{U_{d}}$=0.0 eV) geometry. Like the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (U=$\mathrm{U_{d}}$=0.0 eV) method, HSE06+D3 also fails to localize the defect state due to the absence of the localized geometric distortion near a single Mn atom. Fig. 4C shows the HSE06+D3 DOS of a K-intercalated MnO2 obtained using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (U=$\mathrm{U_{d}}$=1.8 eV) geometry. Here, the defect state appears as in the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (U=$\mathrm{U_{d}}$=1.8 eV) method, but deep in the band gap. Fig. 4D shows the HSE06+D3 DOS of a K-intercalated MnO2 obtained using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (U=1.71 eV and $\mathrm{U_{d}}$= 1.58eV) geometry. Here, also, the defect state appears as in the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (U=$\mathrm{U_{d}}$=1.8 eV) method, but deep in the band gap.
We also calculated HSE06 and HSE06+D3 DOS using an exact exchange mixing parameter $\alpha$ of 0.22 as determined by Peng et al. Peng et al. (2017) on r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry with U=$\mathrm{U_{d}}$=1.8 eV and obtained DOS as shown in Fig. S1 A and B, respectively. As expected, HSE06($\alpha$=0.22) and HSE06($\alpha$=0.22)+D3 DOS do not look different. Notably, the HSE06($\alpha$=0.22) DOS does not look different from the full HSE06($\alpha$=0.25)+D3 evaluated on the same geometry, which is shown in Fig. 4C.
R. Ding et al. Ding et al. (2021) have proposed a position of the polaron close to the CB in layered MnO2 with an alternation of polaron-rich and polaron-poor layers. This scenario matches better with r2SCAN+rVV10+U+$\mathrm{U_{d}}$ DOS, where the polaron is close to the CB.
r2SCAN+rVV10+U+$\mathrm{U_{d}}$ with U=$\mathrm{U_{d}}$=1.8 eV, and U=1.71 eV and $\mathrm{U_{d}}$=1.58 eV give similar DOS. HSE06+D3 DOS are also similar for those geometries, with the polaron appearing deep in the gap.
It is challenging to conclude which method provides a better description of the defect. However, we can assert that all three methods- HSE06, r2SCAN+rVV10+U+$\mathrm{U_{d}}$ and HSE06+D3 using r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry- have successfully described the polaronic defect in birnessite with the proper U and $\mathrm{U_{d}}$ for r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculations.
II.2 Layered NiO2
Our next system under study was a layered NiO2. This material has a hexagonal crystal structure with space group P63/mmc. The Materials Project website (https://next-gen.materialsproject.org/) reports that the material is synthesizable but not stable. Whether stable or not, this material is of interest for our study. Layered NiO2 has a similar structure to birnessite, where Ni ions are in a +4 OS with completely occupied $\mathrm{t_{2g}}$ states, and empty $\mathrm{e_{g}}$ states Martirez and Carter (2019).
We began the DOS calculation using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ functional for both pristine and K-intercalated NiO2 systems. The r2SCAN U value for Ni ions was recently determined by Swathilakshmi et al. Swathilakshmi et al. (2023). For the pristine NiO2, r2SCAN+rVV10+U+$\mathrm{U_{d}}$ with U=$\mathrm{U_{d}}$=2.1 eV, correctly predicts a non-magnetic ground state with a band gap of 1.62 eV, as shown in Fig. 6A. The d-states of Ni(IV) in pristine NiO2 are split into t2g and eg states, with t2g states completely occupied and eg states empty, as shown in Fig. 8A. As in the birnessite case, we expect that adding an extra K atom between layers would transfer an electron from the inserted K atom to a Ni site, forming a defect. The defected Ni site would undergo Jahn-Teller distortion, localizing the electron in the eg state, as shown in Fig. 8B.
Figure 6: r2SCAN+rVV10+U+$\mathrm{U_{d}}$ spin-resolved density of states per atom in (A) pristine NiO2 with Ni(IV) ions with U=$\mathrm{U_{d}}$=2.1 eV and in a single K-intercalated NiO2 with (B) U=$\mathrm{U_{d}}$ =0.0 eV, (C) U=$\mathrm{U_{d}}$=2.1 eV, (D) U= $\mathrm{U_{d}}$ =2.06 eV. The geometry was optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing   5.42 Å. Here we can see one localized occupied state and several localized unoccupied states on the defect Ni(III) cation. The insets in Figs. A and B show the DOS near the 0 eV region. The inset in Fig. B shows that for U=$\mathrm{U_{d}}$=0 the extra electron from the intercalated K goes to the bottom of the conduction band, making a semi-metal.
Figure 7: HSE06+D3 spin-resolved density of states per atom using r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry of Fig. 6 in (A) pristine NiO2 with U=$\mathrm{U_{d}}$=2.1 eV and a single K-intercalated NiO2 with (B) U=$\mathrm{U_{d}}$=0.0 eV and (C) U=$\mathrm{U_{d}}$=2.1 eV, (D) U= $\mathrm{U_{d}}$=2.06 eV. The insets in Figs. A and B show the DOS near the 0 eV region. The inset in Fig. B shows that for U=$\mathrm{U_{d}}$=0 the extra electron from the intercalated K goes to the bottom of the conduction band, making a semi-metal.
Figure 8: d-orbital splitting (A) in Ni(IV) in pristine NiO2 demonstrating filled t2g states and empty eg states and (B) in Ni(III) in a single K-intercalated NiO2
For a K-intercalated NiO2, we initially used U=$\mathrm{U_{d}}$=0.0 eV to observe how r2SCAN+rVV10 without U correction performs for this system, and we obtained the DOS as plotted in Fig. 6B. Again, this choice of U and $\mathrm{U_{d}}$ failed to describe the defect state, as can be seen from the inset in Fig. 6B. The extra electron goes to a delocalized state at the bottom of the conduction band, with fractional occupation on each Ni ion in the supercell, and no individual geometric distortion is observed at any nickel atom. We then used the available U value by setting U=$\mathrm{U_{d}}$=2.1 eV and obtained the DOS, as plotted in Fig. 6C. This choice of U values forms a defect at one of the Ni sites, evidenced by the expected geometric distortion (see Fig. S3 B). The polaron is localized and appears in the gap just below the CB minimum. The formation of the polaron at the VB edge indicates that a K -intercalated NiO2 could show OER catalytic activity similar to that of birnessite by lowering the overpotential. Most of the charge from the K atom is transferred to this defect site, which has a magnetic moment of 0.786 $\mu$B . However, we observe a Ni site in another layer picking up a small but nonzero magnetic moment of 0.137 $\mu$B, with the magnetic moment of all remaining Ni sites smaller than 0.03 $\mu$B. This suggests that we might need different values of U and $\mathrm{U_{d}}$ for r2SCAN+rVV10+U+$\mathrm{U_{d}}$ to accurately describe charge transfer to the defect site in a K-intercalated NiO2 system.
[…]
We also determined the U value of 2.06 eV for the Ni(III) ion by equating the HSE06+D3 magnetic moment to the r2SCAN+rVV10+U magnetic moment, as discussed in the Supplementary Material. As this value is not very different from 2.1 eV, we got a similar r2SCAN+rVV10+U+$\mathrm{U_{d}}$ result with U=$\mathrm{U_{d}}$=2.06 eV compared to U=$\mathrm{U_{d}}$=2.1 eV. Fig. 6D shows r2SCAN+rVV10+U+$\mathrm{U_{d}}$ DOS with U=$\mathrm{U_{d}}$=2.06 eV.
Similar to birnessite, the d-states corresponding to Ni(III) in the VB are shifted to higher energy, as shown in Figs. 6C and 6D, and there is spin symmetry breaking on the total DOS in the CB. Here we do not claim U=0.0 eV to be the precise U value for Ni(IV) ions, but we suspect that U could be different than $\mathrm{U_{d}}$ for a K-intercalated NiO2 system.
We also did HSE06+D3 calculations using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry, and the results are shown in Fig. 7. We observe an increase in the band gap of pristine NiO2, as shown in Fig. 7A. For a single K-intercalated NiO2, due to the absence of a distorted nickel center, we do not observe polaron formation for U=$\mathrm{U_{d}}$=0.0 eV, similar to r2SCAN+rVV10+U+$\mathrm{U_{d}}$ as shown in Fig. 7B; however, we observed an increase in the band gap. We also performed HSE06+D3 calculations for U=$\mathrm{U_{d}}$=2.1 eV and for U=$\mathrm{U_{d}}$=2.06 eV and obtained DOS as shown in Figs. 7C, and 7D, respectively. We observe an increase in the band gap and a shift of the polaron peak deep into the band gap region compared to the corresponding r2SCAN+rVV10+U+$\mathrm{U_{d}}$ DOS. We also observe other effects like shifting of Ni(III) d-states to higher energy and spin symmetry breaking on the total DOS in the CB.
Both methods, full r2SCAN+rVV10+U+$\mathrm{U_{d}}$ and HSE06+D3 using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry with proper U and $\mathrm{U_{d}}$ values, successfully describe defects in a single K-intercalated NiO2. Among the methods employed to study defect in a K-intercalated NiO2, full r2SCAN+rVV10+U+$\mathrm{U_{d}}$ is notably superfast.
II.3 Layered $\mathrm{KCoO_{2}}$
We studied one final related layered material: KCoO2, which has a hexagonal crystal structure with space group P63/mmc, similar to the above materials MnO2 and NiO2. The Materials Project website (https://next-gen.materialsproject.org/) shows a similar material, LiCoO2, with space group P63mc, to be synthesizable but not stable. The electronic configuration of cobalt d-states is distinct from that of Mn and Ni, as illustrated in Fig. 1 in Ref. Ding et al. (2021). The optimal U value determined for Co ions for r2SCAN is 1.8 eV Swathilakshmi et al. (2023).
First, we calculated the DOS of KCoO2 using the optimal U value of 1.8 eV, as shown in Fig. 9A. We observed symmetry in the spin resolved total DOS and Co d-states projected onto a Co site. This suggests that Co(III) ions in KCoO2 have zero spin. The zero spin of Co(III) is also reported in the work of Chen et al. Chen et al. (2011). The d-state configuration of Co(III) is as shown in Fig. 11A.
The pristine CoO2 has cobalt d states in a d5 configuration, and KCoO2 has cobalt in a d6 configuration. A K-intercalated $\mathrm{K_{1.03}CoO_{2}}$ would have a defect cobalt site in a d7 configuration, with the remaining Co ions in a d6 configuration. It is interesting to note that the Co(III) ion has no unpaired electron, but Co(II) has three unpaired electrons due to a low-spin to high-spin crossover Sun et al. (2022); Chen et al. (2011); Bunting et al. (2018); Kim et al. (2023). So, unlike one electron in $\mathrm{e_{g}}$, the state of defected Ni(III) and Mn(III), defected Co(II) has two electrons in the $\mathrm{e_{g}}$ states, with d-state configuration $\mathrm{t_{2g}^{5}e_{g}^{2}}$ (Figure 11 B, C) Sun et al. (2022).
Figure 9: r2SCAN+rVV10+U spin-resolved density of states per atom in (A) pristine KCoO2 with Co(III) ions with U=$\mathrm{U_{d}}$=1.8 eV, a single K-intercalated KCoO2 with (B) U=$\mathrm{U_{d}}$=0.0 eV, (C) U=$\mathrm{U_{d}}$=1.8 eV, (D) U=$\mathrm{U_{d}}$=0.9 eV. The more complicated pattern of in-gap states could reflect a Co(II) configuration different from a single unpaired electron in an eg state. Here we see the appearance of several localized occupied states on the defect Co(II) cation. The geometry was optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing   5.98 Å. Like the d-states of Ni(IV), Co(III) has no electron in the eg state and completely filled t2g states. But, unlike the d-state of Ni(III), which has only one electron in the eg state and completely filled t2g states, Co(II) has four electrons in eg states, two in each channel, and partially filled t2g states, as shown in Fig. 11B. The DOS peaks appearing in up and down channels in the gap region of a K-intercalated KCoO2 are due to the eg electrons in the up and down channels of Co(II), respectively.
Figure 10: HSE06+D3 spin-resolved density of states per atom using r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry of Fig. 9 in (A) pristine KCoO2 with U=$\mathrm{U_{d}}$=1.8 eV, and in a single K-intercalated KCoO2 with (B) U=$\mathrm{U_{d}}$=0.0 eV, (C) U=$\mathrm{U_{d}}$=1.8 eV, and (D) U=$\mathrm{U_{d}}$=0.9 eV.
Figure 11: Schematic diagram of energy of d-orbitals in (A) Co(III) in pristine KCoO2 based on HSE06+D3 calculation using r2SCAN +rVV10+U+$\mathrm{U_{d}}$ geometry with U=$\mathrm{U_{d}}$=1.8 eV, (B) Co(II) in a single K-intercalated KCoO2 based on orbital energy from r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculation with U=$\mathrm{U_{d}}$=1.8 eV and (C) Co(II) in a single K-intercalated KCoO2 based on orbital energies from the HSE06+D3 calculation using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry with U=$\mathrm{U_{d}}$=1.8 eV.
We performed r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculations to study the defects in a K- intercalated KCoO2. First, we tried U=$\mathrm{U_{d}}$=0.0 eV to see how r2SCAN+rVV10 performs without U correction, and obtained a DOS as shown in Fig. 9B. The blue plot, which is the projected density of states (PDOS) of d-states of the Co(II) ion, shows a significant change in the d-state of Co(II) compared to the d-state of Co(III) ion. Our calculation shows the magnetic moment of the Co(II) ion around 3 , indicating three unpaired electrons in d-states. The presence of three unpaired electrons was previously seen in Refs. Sun et al. (2022); Chen et al. (2011); Kim et al. (2023); Bunting et al. (2018). We observed Co(II) d-state splitting, as shown in Fig. 11B. This explains the appearance of Co(II) d-states just above the VB maximum in both channels, absence of Co(II) d-states in the CB in the up channel and presence in the down channel. There is a breaking of spin symmetry in the DOS, due to the appearance of a spin moment on the defected Co ion. The phenomena like shifting of d-states to higher energy and spin symmetry breaking in CB are similar to those observed and explained before in Peng et al.’s work Peng et al. (2017) in the study of birnessite.
Then we used U=$\mathrm{U_{d}}$=1.8 eV in r2SCAN+rVV10+U+$\mathrm{U_{d}}$ and obtained a DOS as shown in Fig. 9C. We do not see much difference between the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ DOS for U=$\mathrm{U_{d}}$=0.0 eV and U=$\mathrm{U_{d}}$=1.8 eV. This indicates that the occupied Kohn-Sham orbitals are well localized and there is no fractional occupancy in r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculations with U=$\mathrm{U_{d}}$=0.0 eV. However, the larger U appears to bring peaks of the defected Co(II) d-states closer together in the band gap region just above the VB maximum.
As discussed in the Supplementary Material, we tried to determine the U value for the Co(II) ion by equating the HSE06+D3 magnetic moment to the r2SCAN+rVV10+U magnetic moment and obtained the U value for the Co(II) ion to be 0.91 eV. We used U=$\mathrm{U_{d}}$=0.91 eV and obtained a DOS as shown in Fig. 9D, which looks similar to Figs. 9B and 9C.
We tried to compare a full HSE06+D3 DOS with r2SCAN+rVV10+U+$\mathrm{U_{d}}$ for pristine KCoO2 and a K-intercalated KCoO2. We obtained the KCoO2 supercell by inserting a K layer for each CoO2 layer. Compared to the pristine CoO2 supercell, which contains 96 ions, the KCoO2 supercell has 32 extra K ions, and the K-intercalated KCoO2 supercell has 33 extra K ions. Therefore, one would expect a different lattice parameter for KCoO2 and a K-intercalated KCoO2 , which requires the full relaxation of the supercell. We attempted to optimize the structure using the HSE06+D3 method, but it is very challenging for these systems in terms of time and resource. After a month, we chose not to proceed further. This challange highlights one of the important reasons that motivated us to explore the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method, highlighting the significance of the r2SCAN meta-GGA for defect studies.
However, we calculated the HSE06+D3 DOS using r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry. As shown in Fig. 10A, HSE06+D3 predicts KCoO2 to be non- magnetic with an increased band gap of 4.17 eV. For a K-intercalated system with r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry, the HSE06+D3 DOS is shown in Figs. 10 B, C, and D for U=$\mathrm{U_{d}}$=0.0 eV, 1.8 eV, and 0.91 eV, respectively. Here we see that HSE06+D3 not only opens the gap between VB and CB, but it also brings the peaks of the defected Co(II) d-states closer together in the band gap region just above the VB maximum. We observed Co(II) d-states splitting as shown in Fig. 11C for the HSE06+D3 calculation using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry with U=$\mathrm{U_{d}}$=1.8 eV.
Fig. 11 shows our electronic configurations for Co(III) and Co(II) in octahedral or nearly octahedral coordination with oxygen. For Co(II), we have seven d electrons. There are three $\mathrm{t_{2g}}$ orbitals of each spin, and two $\mathrm{e_{g}}$ orbitals of each spin, and we can put no more than one electron into each spin orbital. Our output tells us the relative energy order of each occupied d state of each spin, that the net spin is nearly that of three spin-up electrons, and that the lowest-energy orbitals are spin-up while the highest-occupied orbital is spin-down, but not whether a given spin orbital is $\mathrm{t_{2g}}$ or $\mathrm{e_{g}}$. So we begin by putting three spin-up electrons into $\mathrm{t_{2g}}$ orbitals, to minimize the electronic repulsion. We assume that the $\mathrm{e_{g}}$ electrons prefer to be spin-up, to take advantage of an exchange interaction with the three up-spin $\mathrm{t_{2g}}$ orbitals, and that spin-up $\mathrm{e_{g}}$ levels will compete in energy with the spin-down $\mathrm{t_{2g}}$ levels. Then the configuration will be the one shown in Fig. 11 B and C, $\mathrm{(t_{2g}\uparrow)^{3}(t_{2g}\downarrow)^{2}(e_{g}\uparrow)^{1}(e_{g}\uparrow)^{1}}$. This result agrees with the $\mathrm{(t_{2g})^{5}(e_{g})^{2}}$ configuration found for Co(II) in $\mathrm{Li_{2}CoO_{2}}$ in Ref. Sun et al. (2022). This electronic configuration is also supported by the total charge density plots (see Fig. S5) of the four localized orbitals that appear in the band gap region of $\mathrm{K_{1.03}CoO_{2}}$.
Regardless of coordination, Co(II) has been found to have a net spin nearly equal to that of three spin-up electrons in Refs. Chen et al. (2011); Sun et al. (2022); Bunting et al. (2018); Kim et al. (2023), but the distribution over $\mathrm{t_{2g}}$ and $\mathrm{e_{g}}$ of course depends upon coordination and other details of the nearby environment.
Fig. S4 B of the Supplementary Material shows the Jahn-Teller distorted tetrahedron around a defect Co(II) ion, with two of the Co(II)-O bonds about 10% longer than the remaining four. The angles are also distorted.
III Conclusions
Defects in solids can be utilized to tune the electronic structure of materials, impacting applications in various fields. Defects can be investigated using first-principles calculations with DFT. However, DFAs introduce SIE that could vary depending on the element, system, site, and oxidation state, hindering the performance of DFAs. Hybrid functionals mitigate SIE by mixing exact exchange , but at the expense of computational cost, successfully describing defects in many materials. The computational cost of hybrid functionals increases significantly for a SCF (self-consistent field) calculation compared to meta-GGAs like r2SCAN. Defect studies require large supercells, making SCF calculations very expensive. Ionic relaxation calculations for defect studies involve many SCF cycles, making the process cumbersome for a hybrid functional to optimize the structure.
Here we demonstrate that the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method can successfully describe defects in materials with proper values of U and $\mathrm{U_{d}}$. The method is much faster than hybrid functionals but equally accurate. Using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometry enables the completion of hybrid functional calculations in a reasonable time and with fewer resources. This approach makes hybrid functional calculations feasible for larger systems with transition-metal ions.
The r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method proposed here is not only faster but also as accurate as the already established hybrid functional method for defect studies. This method can expedite the study of defects in materials. The r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method can also be employed in systems without defects where ions of the same species are in different OS.
We have studied the layered oxides MnO2, NiO2, and KCoO2, both in the pristine state and with one additional K atom intercalated between layers in a supercell. Inexpensive r2SCAN+rVV10+U+$\mathrm{U_{d}}$ equilibrium geometries have been used for electronic structure calculations with r2SCAN+rVV10+U+$\mathrm{U_{d}}$ and with the expensive HSE06 hybrid functional. For K-intercalated MnO2 and NiO2, we find no localized eg state on the defected transition metal ion for U=$\mathrm{U_{d}}$=0, but we find one such state for standard positive U values. This state, in the gap between conduction and valence bands, accepts the electron donated by the intercalated K. (For U=$\mathrm{U_{d}}$=0, the extra electron from the intercalated K goes into the bottom of the conduction band, making the intercalated material a semi-metal. The extra electron is then delocalized over the supercell, with fractional occupation on each transition-metal ion.)
For K-intercalated KCoO2, both for U=$\mathrm{U_{d}}$=0 (standard semilocal r2SCAN without a nonlocal +U self-interaction correction) and for standard positive U values, we seem to find two majority-spin and two minority-spin occupied localized eg states in the gap between valence and conduction bands, and three empty localized minority-spin t2g states in the gap above the conduction band. This surprising result is consistent with a dramatic change in the electronic configuration (Fig. 11) from the undefected Co(III) ions to the defected Co(II) ion, as reported previously in Ref. Chen et al. (2011).
The r2SCAN+rVV10+U+$\mathrm{U_{d}}$ approach is successful, but it has limitations. It requires material-dependent parameters U, and those are typically available for only a few oxidation states of each transition metal. We have suggested a way to supplement the available values by choosing U to match the HSE06 magnetic moments, but that does not work when a moment is zero. A possibly more satisfactory approach would be to do r2SCAN+rVV10+U calculations for isolated transition-metal ions with integer M and half-integer M+1/2 electron numbers. Since the +U correction should be much less important for integer than for half-integer electron number, one can use U(M+1/2) for both M and M+1. Then U(M+1/2) could be found to satisfy the PPLB Perdew et al. (1982) condition E(M+1/2) = [E(M+1)+E(M)]/2, and U($\mathrm{M^{\prime}}$) could be found for all $\mathrm{M^{\prime}}$ by interpolation of U($\mathrm{M^{\prime}}$+1/2), yielding a set of oxidation-state-dependent values for each transition element. This possibility, which requires further thought and computation, is under consideration now in our group. Useful transfer of U from the free ion to the ion in a solid is not guaranteed, but the ratios of the U values for one element in different oxidation states might transfer better.
IV Computational Details
First-principles calculations were performed with the projector-augmented wave method Blöchl (1994), implemented in the VASP code Kresse and Hafner (1994); Kresse and Joubert (1999). A 4 x 4 x 1 supercell was used to simulate defects in layered TMOs. For all supercell calculations, a 2 x 2 x 2 $\Gamma$-centered Monkhorst-Pack k mesh Monkhorst and Pack (1976) was used.
For r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculations, a 500 eV cutoff for the plane waves was used. In all r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculations, the cell volume was relaxed with the ISIF=3 setting until forces converged to less than 0.03 eV/Å and energy converged to less than 10-6 eV. To conduct r2SCAN+rVV10+U+$\mathrm{U_{d}}$ calculations, we employed the simplified rotationally invariant framework developed by Dudarev et al. Dudarev et al. (1998).
For HSE (as for r2SCAN+rVV10+U+$\mathrm{U_{d}}$), corrections due to the periodic boundary condition and the supercell size have been added Peng et al. (2017); Kumagai and Oba (2014). For Fig. 2, we used an exact mixing parameter $\alpha$ of 0.22 in the hybrid functional HSE06 Peng et al. (2017); Krukau et al. (2006b), as determined from the Generalized Koopman’s Condition (GKC) method Lany and Zunger (2009); Lany (2011); Peng and Lany (2012); Peng et al. (2015). For all the remaining HSE06+D3 calculations, we used the standard exact exchange mixing parameter of 0.25. A plane-wave basis with an energy cutoff of 400 eV was employed.
V Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science Basic Energy Sciences under Award No. DE-SC0023356.
VI Supplementary Material
Determining a U parameter from the HSE06 magnetic moment. Effect of the D3 dispersion correction on lattice constants and density of states. Jahn-Teller symmetry breaking of the CoO6 tetrahedron in a single K-intercalated KCoO2.
Supplementary Material
Determination of Hubbard U from Magnetic Moment
To determine a Hubbard U for some of our calculations, we equated the HSE06+D3 magnetic moment (MM) of ions inside the Wigner-Seitz sphere with the r2SCAN+rVV10+U’s MM. First, we relaxed the structure using the HSE06+D3 method and used this structure to calculate the HSE06+D3 MM. We then used the HSE06+D3 structure to do the r2SCAN+rVV10+U calculation, where we slowly varied the U value until r2SCAN+rVV10+U’s MM becomes equal to the HSE06+D3’s MM. We used smaller bilayer unit cells of transition metal oxides for U determination using this method. We used MnO2, KMnO2, KNiO2, and CaCoO2 to get U values for Mn(IV), Mn(III), Ni(III), and Co(II) ions, respectively. All calculations are performed using the VASP code, where we employed a Gamma-centered Monkhorst-pack grid of size 8 x 8 x 8 and a cut off energy of 400 eV for the plane-wave basis.
First, we used HSE06+D3 with an exact exchange mixing parameter ($\alpha$) of 0.22 to be consistent with Peng et. al.’s work Peng et al. (2017). We have summarized the calculated U value for these ions in Table S1 below.
HSE06($\alpha=0.22$)+D3 MM ($\mu_{B}$) r2SCAN+rVV10+U MM ($\mu_{B}$) U (eV)
MnO2 3.023 3.023 1.41
KMnO2 3.933 3.933 1.26
KNiO2 1.098 1.098 1.53
CaCoO2 2.663 2.663 0.55
Table S1: The U value determined by equating the magnetic moments (MM) of HSE06+D3 ($\alpha$ = 0.22) and r2SCAN+rVV10+U, using the HSE06+D3 ($\alpha$ = 0.22) geometry for both functionals.
We could not use this method to determine the U values for Ni(IV) and Co(III) ions using NiO2 and KCoO2 systems because the MM of Ni(IV) and Co(III) ions in NiO2 and KCoO2 is zero, largely independent of the U value, in the r2SCAN+rVV10+U calculation. So, we used the same U values for Ni(IV) and Co(III) as determined for Ni(III) and Co+4, respectively. We used the U values determined in Table S1, as well as U values from the literature Ding et al. (2021), and calculated lattice parameters for MnO2, NiO2, and CaCoO2 unit cells to see how r2SCAN+rVV10+U performs compared to HSE06+D3 with $\alpha$=0.22. The results are summarized in Table S2 below.
  System Method Lattice Parameters (Å) Lattice Angles (Degree)
a b c $\alpha$ $\beta$ $\gamma$
r2SCAN+rVV10 2.82 2.82 9.23 90 90 120
MnO2 r2SCAN+rVV10+U(1.41 eV) 2.83 2.83 9.22 90 90 120
HSE06($\alpha=0.22$)+D3 2.83 2.83 9.59 90 90 120
r2SCAN+rVV10 2.74 2.74 8.98 90 90 120
NiO2 r2SCAN+rVV10+U(1.53 eV) 2.73 2.73 9.00 90 90 120
HSE06($\alpha=0.22$)+D3 2.72 2.72 9.25 90 90 120
r2SCAN+rVV10 2.89 2.89 11.64 90 90 120
KCoO2 r2SCAN+rVV10+U(0.55 eV) 2.89 2.89 11.65 90 90 120
HSE06($\alpha=0.22$)+D3 2.86 2.86 11.21 90 90 120
Table S2: Lattice parameters and lattice angles comparison for r2SCAN+rVV10, r2SCAN+rVV10+U and HSE06($\alpha=0.22$)+D3 methods for the systems MnO2, NiO2 and KCoO2 .
We see that the lattice constants agree well for r2SCAN+rVV10 and r2SCAN+rVV10+U, but they disagree with HSE06+D3 ($\alpha$=0.22) lattice parameters, especially the c parameter. Probably this is because the D3 parameters were determined for HSE06 at $\alpha$ = 0.25, not 0.22.This motivated us to go back to full HSE06.
HSE06($\alpha=0.25$)+D3 MM ($\mu_{B}$) r2SCAN+rVV10+U MM ($\mu_{B}$) U (eV)
MnO2 3.036 3.036 1.71
KMnO2 3.947 3.947 1.58
KNiO2 1.163 1.163 2.06
CaCoO2 2.683 2.683 0.91
Table S3: The U value determined by equating the magnetic moments (MM) of HSE06+D3 ($\alpha$ = 0.25) and r2SCAN+rVV10+U, using the HSE06+D3 ($\alpha$ = 0.25) geometry for both functionals.
As above, we again determined U values by comparing the MM of r2SCAN+rVV10+U with the full HSE06+D3 ($\alpha$=0.25 ) functional. Table S3 below summarizes the result. The U values determined in the Table S3 are closer to the values in the literature, where optimal values for Mn, Ni, and Co ions for r2SCAN are 1.8 eV, 2.1 eV, and 1.8 eV, respectively Ding et al. (2021). The difference in oxidation states of ions could be the cause of the discrepancy. We used the above U values for the r2SCAN+rVV10+U functional and again compared the lattice parameters. The results are summarized in Table S4 below.
  System Method Lattice Parameters (Å) Lattice Angles (Degree)
a b c $\alpha$ $\beta$ $\gamma$
r2SCAN+rVV10 2.82 2.82 9.23 90 90 120
MnO2 r2SCAN+rVV10+U(1.71 eV) 2.84 2.84 9.11 90 90 120
HSE06($\alpha=0.25$)+D3 2.82 2.82 9.25 90 90 120
r2SCAN+rVV10 2.74 2.74 8.98 90 90 120
NiO2 r2SCAN+rVV10+U(2.06 eV) 2.73 2.73 8.98 90 90 120
HSE06($\alpha=0.25$)+D3 2.73 2.73 9.03 90 90 120
r2SCAN+rVV10 2.89 2.89 11.64 90 90 120
KCoO2 r2SCAN+rVV10+U(0.91 eV) 2.89 2.89 11.65 90 90 120
HSE06($\alpha=0.25$)+D3 2.87 2.87 11.61 90 90 120
Table S4: Lattice parameters and lattice angles comparison for r2SCAN+rVV10, r2SCAN+rVV10+U and HSE06($\alpha=0.25$)+D3 methods for the systems MnO2, NiO2 and KCoO2.
Figure S1: Spin-resolved density of states per atom using r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (with U=$\mathrm{U_{d}}$=1.8 eV) geometry in a single K-intercalated MnO2 using (A) HSE06 (with $\alpha$=0.22) and (B) HSE06 (with $\alpha$=0.22)+D3.
Figure S2: (A) Mn(IV) - O bond lengths in pristine MnO2 and (B) defect Mn(III) - O bond lengths in a K-intercalated MnO2 ($\mathrm{K_{0.03}MnO_{2}}$) calculated using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (with U=$\mathrm{U_{d}}$=1.8 eV) functional, showing the Jahn-Teller symmetry breaking of the MnO6 octahedron. All lengths are in angstrom units. The red balls represent O ions and purple Mn ions. In the pristine materials, the three $\mathrm{t_{2g}\ (d_{xy},d_{yz},d_{zx})}$ space orbitals are equally populated, and the two $\mathrm{e_{g}\ (d_{x^{2}-y^{2}},d_{z^{2}})}$ spatial orbitals (whose lobes point from the metal cation to the 6 oxygens on the corners of the octahedron) are empty.
Figure S3: (A) Ni(IV) - O bond lengths in pristine NiO2 and (B) defect Ni(III) - O bond lengths in a K-intercalated MnO2 ($\mathrm{K_{0.03}NiO_{2}}$) calculated using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (with U=$\mathrm{U_{d}}$=2.1 eV) functional, showing the Jahn-Teller symmetry breaking of the NiO6 octahedron. All lengths are in angstrom units. The red balls represent O ions and silver Ni ions. In the pristine materials, the three $\mathrm{t_{2g}\ (d_{xy},d_{yz},d_{zx})}$ space orbitals are equally populated, and the two $\mathrm{e_{g}\ (d_{x^{2}-y^{2}},d_{z^{2}})}$ spatial orbitals (whose lobes point from the metal cation to the 6 oxygens on the corners of the octahedron) are empty.
Figure S4: (A) Co(III) - O bond lengths in pristine KCoO2 and (B) defect Co(II) - O bond lengths in a K-intercalated KCoO2 ($\mathrm{K_{1.03}CoO_{2}}$) calculated using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ (with U=$\mathrm{U_{d}}$=1.8 eV) functional, showing the Jahn-Teller symmetry breaking of the CoO6 octahedron. All lengths are in angstrom units. The red balls represent O ions and blue Co ions. In the pristine materials, the three $\mathrm{t_{2g}\ (d_{xy},d_{yz},d_{zx})}$ space orbitals are equally populated, and the two $\mathrm{e_{g}\ (d_{x^{2}-y^{2}},d_{z^{2}})}$ spatial orbitals (whose lobes point from the metal cation to the 6 oxygens on the corners of the octahedron) are empty.
Figure S5: $\mathrm{r^{2}SCAN+rVV10+U+U_{d}\ (U=U_{d}=1.8\ eV)}$ based total charge density plots of four localized orbitals in the band gap region of $\mathrm{K_{1.03}CoO_{2}}$ around the defect Co(II) ion . The plots in the upper panel correspond to the two highest occupied orbitals in the spin up channel and clearly show them to be the $\mathrm{e_{g}}$ orbitals. The lower panel corresponds to the two highest occupied orbitals in the spin down channel and shows them to be the $\mathrm{t_{2g}}$ orbitals.